Circular No. 8424

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2004 H. E. Bond, Space Telescope Science Institute; F. Walter, Stony Brook

University; and J. Espinoza, D. Gonzalez, and A. Pasten, SMARTS Consortium, Cerro Tololo Interamerican Observatory, report spectroscopic and CCD observations with the 1.5-m and 1.3-m Cerro Tololo telescopes of the object discovered by Liller (IAUC 8422). Spectra obtained on Oct. 21.30, 22.33, 23.28, 24.31, and 25.44 UT (resolution 0.43 nm and range 353–530 nm, except resolution 0.16 nm and 387–454 nm on Oct. 23.28, and resolution 0.43 and range 600–780 on Oct. 25.44) show strong, triple-peaked Balmer emission lines, as well as He I and strong N III emission. The Balmer profiles consist of a sharp core redshifted by 0.4 nm (consistent with the radial velocity of the LMC) atop a broad, double-horned pedestal. The two outer peaks forming the horned structure have velocities of ∼ ±2500 km/s. A blueshifted wind absorption feature, whose velocity on Oct. 21.30 was −3700 km/s, has gradually decreased in strength. On Oct. 25.33, the Hα line had FWZI = 9800 km/s, with an equivalent width of 120 nm. Astrometry of the CCD frames using USNO-B1.0 reference stars gives α = 5^{h}56^{m}42^{s}.42, δ = −68^{o}54′34′′.8 (equinox 2000.0); the faint star mentioned on IAUC 8422 has USNO-B1.0 end figures 42^{s}.40, 34′′.6, suggesting that it is likely the quiescent counterpart. A comparison by the undersigned of a B-band Cerro Tololo image, supplied by Bond, with the Harvard/Boyden 0.61-m Bruce telescope plate A19851 (a 45-min exposure taken on the night of 1937 Nov. 23–24 in poor seeing) shows that the distorted 1937 nova (YY Dor) image coincides in position with the 2004 nova, though the centroid of the 1937 photographic image may be a bit to the southwest of the centroid of the 2004 CCD image. (YY Dor would then be only the second recurrent nova known in the LMC.) Photometry of the 2004 nova from the CCD frames: Oct. 21.33, V = 12.15, B − V = −0.14; 23.24, 12.79, −0.25; 24.24, 13.00, −0.30; 25.20, 13.16, −0.23.

E. Mason, A. Ederoclite, M. Stefanon, and T. H. Dall, European Southern Observatory, Santiago; and M. Della Valle, Istituto Nazionale di Astrofisica, Arcetri/Firenze, report that a spectrum of the new nova, obtained on Oct. 25.3 UT with the La Silla 2.2-m telescope (+ Fiber-fed Extended Range Optical Spectrograph; range 390–900 nm, resolution 48000), is dominated by strong Balmer lines and O I (4) at 844.6 nm. Emission from He I (11 and 10) was also detected. The line profiles are very broad. with an average FWHM of 6600 km/s; such a large velocity is consistent with the typical spectra of recurrent novae at maximum.

2004 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green